OpenStudy (anonymous):

A catapult launches a boulder with an upward velocity of 112 ft/s. The height of the boulder, h, in feet after t seconds is given by the function h=-16t^2+112t+30 How long does it take the boulder to reach its maximum height? What is the boulder’s maximum height? Round to the nearest hundredth, if necessary.

OpenStudy (primeralph):

Generally, Max Height is at x = -b/(2a) = 112/(32) seconds.

OpenStudy (primeralph):

@Kenziehulse You have to tell me whether or not you understand.

OpenStudy (anonymous):

No i kinda dont understand.

OpenStudy (primeralph):



OpenStudy (primeralph):